Hello! How can I find the focus and the directrix of the following equation: (y+2)^2=-12(x-5)
Thank you!

Answer by rothauserc(4718)

(y+2)^2 = -12(x-5)
:
This is an equation of a parabola in standard form, this parabola opens to the left since p is negative
:
The standard form is
:
( y - k )^2 = 4 * p * ( x - h )
:
The vertex is (h, k) = (5, -2)
:
Note that p = -12 / 4 = -3
:
Focus = (h + p, k) = (5-3, -2) = (2, -2)
:
Directrix is x = h - p = 5 - (-3) = 8
: